# SimLOB: Learning Representations of Limited Order Book for Financial Market Simulation

### v3

@@ Abstract
1 Introduction
One important and challenging issue in mining financial market data is to discover the causes of the abnormal market phenomena, e.g., flash crash [1, 2] and spoofing [3, 4]. Traditional machine learning methods are well-equipped to detect the anomalies but lack of the financial interpretability of the causes [5]. Given observed market time series data, financial market simulation (FMS) aims to directly approximate its underlying data generating process, informed by the trading rules of the real market [6]. By simulating the underlying trading activities behind the observed market data, FMS is able to discover how the structure of the financial market changed over time and explain the changes on the micro level [7, 8].
@@ 1 Introduction
Advanced methods from both the fields of optimization and statistical inference have been developed to address the calibration problem [13, 14]. Optimization methods primarily focus on minimizing the discrepancy through black-box optimization techniques [15], while statistical inference approaches estimate the likelihood or posterior distribution of $\hat{\mathbf{X}}_{T}$ and $\mathbf{w}$ for selected samples subject to $D(\hat{\mathbf{X}}_{T},M(\mathbf{w}))<\epsilon$ [16, 17]. In both ways, the calibration is conducted in an iterative randomized sampling manner in the parameter space and the sampling is mostly guided by the discrepancy. Intuitively, the more information observed from the market included in $\hat{\mathbf{X}}_{T}$, the closer $M(\mathbf{w})$ is expected to approximate the underlying data generative process. Most FMS works merely consider the mid-price data as $\hat{\mathbf{X}}_{T}$ [14, 18], other important information of observable data, e.g., trading volume, bid/ask directions, and inter-arrival time of orders. Consequently, the calibrated $M(\mathbf{w})$ fails to capture the underlying market structure.
Compared to mid-price data, the Limit Order Book data (LOB) provides a more detailed and informative source of market information. It offers a comprehensive view of all untraded orders, capturing critical market dynamics beyond simple price changes [19]. As the fundamental market data, LOB is widely adopted by most world-class securities exchanges [20]. In LOB, all the untraded orders fall into either the ask side (for selling) or the bid side (for buying) according to their order directions. When a new order arrives, the LOB is updated immediately in one of the two ways: if the order price matches a price on the opposite side, the order is executed and the corresponding orders on the opposite side are removed; otherwise, the new order is added to its respective side. On each side, the untraded orders are organized in the "price-first-time-second" manner [21]. Notably, the untraded orders in LOB implicitly reflect the trading intentions of the entire market. Besides, the immediate updates to the LOB provide a highly fine-grained view of market microstructure. On this basis, simulating LOB offers a conceptually close approximation to the real data generative process underlying the market scenario of interest.
@@
2 Related Work
2.1 Calibration of FMS
@@ 2.2 Deep Learning for LOB
Recent works propose to generate LOB data with deep learning, which is quite close to the goal of FMS [32]. However, they do not involve any validation or calibration steps to force the generated data resembling any given observed LOB but are only limited to following some macro properties of the market [33, 34, 35]. Besides, those deep generative models lack the interpretability of the quotes and trades in the real market. Some of the models also are not informed by the real-world matchmaking rules for updating LOB [36, 37], and all of them are not explainable in why the specific orders are submitted. These issues impose significant restrictions on the use of these models as the simulators since one can hardly intervene in the order streams to do a "what-if" test and reveal the micro-structured causes of certain financial events [38, 39]. Furthermore, they explicitly require the ground truth order streams as input to train the networks, which is not available in the setting of FMS calibration problems.
@@
3 The Proposals
3.1 The Problem Definition and Challenges
@@
    Data is with different scales. Prices typically fluctuate around several thousands of cents with a tick size of 1 cent, while volumes range from $\xi$ shares to millions of shares, with $\xi$ (e.g., 100 or 1000) varying across exchanges.
To summarize, learning representations of the LOB requires jointly tackling the aforementioned three challenges. While the challenges of multi-scale and non-linearly auto-correlation have been extensively investigated in the literature, the precedence constraints have been rarely seen in other fields. Further more, it is uncommon to address them simultaneously with one representation model. These difficulties jointly offer the LOB data unique challenges and a good arena for studying representation learning of such type of data, particularly as the reconstructed multi-dimensionality data must maintain precedence constraints and avoid "intersection" at each time step.
@@ 3.2 The Network Architecture
Implementation Details. The proposed autoencoder, SimLOB, processes observed LOB by spliting it into multiple segments of $\tau=100$ time steps, as suggested by [30]. Through sensitive analysis presented in Section 4.6, we set $L=2$ and $\tilde{\tau}=128$ by default. The reconstruction error is defined as:
@@
4 Empirical Studies
The following four research questions (RQs) are major concerns in this work.
RQ-1: Can various LOBs be effectively represented as vectors?
RQ-2: Is it really beneficial to calibrate FMS with LOB instead of traditional mid-price?
RQ-3: Does "the better representation of LOB, the better calibration of FMS" generally hold?
RQ-4: What is the recommended architecture?
Four groups of experiments are conducted accordingly.
Group-1: The proposed SimLOB and 8 SOTA networks are trained on a large volume of synthetic LOB. Reconstruction errors on synthetic and real testing data are measured to evaluate vectorized representations.
Group-2: Calibration errors are compared over the representations learned by SimLOB, the raw LOB, and the mid-price.
Group-3: Latent vectors from the above 9 networks are applied to 10 FMS calibration tasks to show the consistency between the reconstruction performance and the calibration errors.
Group-4: Parameter sensitivity and structural ablation of SimLOB are analyzed in terms of reconstruction and calibration.
Figure 3: The workflow of the PGPS financial simulation model.
4.1 General Experimental Setup
The Adopted FMS Model. The widely studied Preis-Golke-Paul-Schneid (PGPS) model [46] is adopted as the FMS model, simulating two types of agents: 125 liquidity providers submiting limit orders and 125 liquidity takers submiting market orders or cancel untraded limit orders. The simulation flowchart is shown in Fig. 3. At each time step $t$-th, a liquidity provider submits a limit order with probability $\alpha=0.5$. A liquidity taker submits a market order with probability $\mu$ and cancel an untraded limit order with probability $\delta$. The probability of the market order being either the bid side or ask side is $q_{taker}(t)$ and $1-q_{taker}(t)$, respectively, where $q_{taker}(t)$ follows a mean-reverting random walk centered at 0.5, with increments of ±$\pm\Delta s$ and reversion probability $0.5+|q_{taker}(t)-0.5|$ Each order has a fixed volume of 100 shares. The price of an ask limit order is determined by $p^{a}_{1}(t)-\lfloor-\lambda(t)\log u)\rfloor-1$ and a bid limit order is determined by $p^{b}_{1}(t)-\lfloor-\lambda(t)\log u)\rfloor+1$, where $\lambda(t)=\lambda_{0}(1+\frac{|q_{taker}(t)-0.5|}{\sqrt{\langle q_{taker}-0.5\rangle^{2}}}C_{\lambda})$. Here $\langle q_{taker}-0.5\rangle^{2}$ indicates a pre-computed value obtained by taking the average of $10^{5}$ Monte Carlo iterations of $(q_{taker}-0.5)^{2}$ before simulation. $u\sim U(0,1)$ is a uniform random number. The price of the market order is automatically set to the best price level on the opposite side. In summary, the PGPS model contains 250 agents and the key parameters to be calibrated are $\mathbf{w}=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$.
Table 1: Parameter Ranges for Generating Synthetic Data
@@ 4.1 General Experimental Setup
The Dataset for Representation Learning. Augmenting training data with the synthetic LOB generated by FMS is increasingly popular [32, 37]. Hence, this work trains the SimLOB with purely the synthetic LOB data generated by the PGPS model with different parameters. In general, each setting of the 6-parameter tuple of PGPS actually defines a specific market scenario. To ensure the synthetic data enjoys enough diversity, we uniformly randomly sample 2000 different settings of those 6-parameter tuples from a predefined range suggested by [26] (see Table 1). Each configuration simulates over 50,000 time steps, yielding 500 sequences with 100 time steps. This results in 1 million LOB sequences ($\tau=100$), with 80% for training and 20% for testing. Additionally, real market data from the sz.000001 stock of the Chinese market (during May of 2019) is used for evaluation.
Table 2: The Characteristics of the 9 Compared Networks
@@
The Compared Networks. This experiment compares 7 SOTA networks and one enhanced version (TransLOB-L), using their architecture (excluding the last linear layer) as representation learning layers, all configured with $\tilde{\tau}=128$. These networks have been frequently adopted in comparisons in recent deep learning based LOB analysis works [30, 47, 43], primarily focusing on price trend prediction. We follow the conventions of [30] to name the 8 networks as MLP[48], LSTM[48], CNN1[49], CNN2[50], CNN-LSTM[50], DeepLOB[43], TransLOB[41]. As their names suggest, despite the fact that MLP, LSTM, CNN1, and CNN2 adopt single-type networks, the other works mostly employ a Convolution-Recurrent architecture to intuitively first extract the prices precedence and then capture the temporal auto-correlation. TransLOB employs the Transformer architecture to replaces the recurrent network while still keeping the convolution layers. TransLOB-L is an enlarged version of TransLOB to keep the size approximately aligned with SimLOB by using 7 stacked Transformers instead of 2 as in the original TransLOB. All 8 networks are trained under the same protocol as SimLOB, with input formats and sizes follow the original papers. Table 2 summarizes their characteristics: the second and third columns list acceptable input formats of each network, and the last column shows the sizes of the networks. The MLP has the largest size of $2.2\times 10^{7}$ weights, which can be simply considered as the enlarged version of $\mathrm{FCN}_{2}$.
@@
Table 3: Parameter Settings of Generating 10 Synthetic Data with the PGPS Model
@@
10 synthetic LOB data are randomly generated with 10 different settings of 6-parameter tuples, each containing $T=3600$ time steps to simulate approximately 1 hour of market activity. The parameters range for randomly sampling parameters is in the Table 1, and the sampled tuples are listed in Table 3. These data are utilized as the target data. Note that these data instances are challenging that they have a much higher frequency than traditional FMS works who can only calibrate to daily data [26]. And the length of the target data is also at least 10×10$\times$10 × longer than the existing calibration works [14]. The objective function for calibration employs MSE. The calibration objective for mid-prices follows traditional works do, Eq.(2) gives
@@
Figure 4: Reconstruction errors on 0.2M testing instances.
Figure 5: The reconstructed bid prices (10 levels) on 5 synthetic data are visualized together with $\mathbf{Err}_{r}$.
Figure 6: The reconstructed bid prices (10 levels) on 4 synthetic data and 4 real market data are visualized together with $\mathbf{Err}_{r}$.
4.2 Group-1: Learning the Vectorized Representations of LOB
The reconstruction errors on 0.2 million synthetic testing data of all 9 compared algorithms are depicted in Fig. 4. Among them, SimLOB not only achieves the smallest averaged reconstruction error, but performs the most stable performance, with reconstruction errors of SimLOB following a power-law distribution and a mode of 0.003. Almost all the CNN-based networks (including CNN1, CNN2, CNN-LSTM, TransLOB, and TransLOB-L) perform the worst, suggesting that CNN is not effective as expected to deal with LOB data. This is quite contradictory to the trend of existing works on adopting deep learning for analyzing LOB. This phenomenon is later studied in-depth in Section IV-E with more empirical ablations.
The reconstructed data of seven compared networks across five synthetic data demonstrates the superiority of SimLOB. These depicted data is randomly selected from the testing instances based on the principles of covering different fluctuating amplitudes and trends. In Fig. 5, the data reconstructed by SimLOB mostly closely resembles the ground truth, while other networks have tended to smooth the target data, failing to fully preserve the key properties of LOB in their latent vectors. TransLOB and TransLOB-L are excluded from visualization as their reconstructed data appear to be nearly flat, consistent with their poor performance shown in Fig. 4. This indicates that the Transformer architecture alone is not the panacea to the representation learning of LOB. And thus it verifies that the importance of the proposed feature extraction and dimension reduction blocks in SimLOB.
Note that these networks are trained purely on synthetic data and are directly applied to reconstruct real market data in a zero-shot manner. In real market data, volume at each price level are much larger than in the PGPS model. Thus the price value at each level fluctuates much smoothly, as shown in the first column of the last 4 rows in Fig. 6. Since the compared networks tend to smooth the reconstructed data, this comparison in Fig. 6 can also serve as the generalization test to see how SimLOB performs on those essentially smooth data and the consistency on different types of target data.
As shown in the last 4 rows of Fig. 6, the compared networks all fail to reconstruct meaningful details. To be specific, the reconstructed bid price levels from CNN1, CNN-LSTM and DeepLOB tend to be close and indistinguishable. Besides, strong fluctuations that do not exist in the real data have emerged. Note that, we only show the reconstructed data from CNN1, CNN-LSTM and DeepLOB here due to the limited space. And the selected three networks cover different classic structures like CNN and LSTM. Contrarily, SimLOB can still capture the information about price precedence and fluctuations to some extent, though the range of prices deviates. This performance aligns with results on synthetic data, as shown in the first 4 rows of Fig. 6, where SimLOB consistently outperforms others. These findings suggest that SimLOB is robust regardless of fluctuations of the target data and can be further generalized to more diverse LOB with additional treatments, such as training with real data.

### v4

# SimLOB: Learning Representations of Limit Order Book for Financial Market Simulation

@@ Abstract
Index Terms:
limit order book, representation learning, financial market simulation, model calibration.
I Introduction
One important and challenging issue in mining financial market data is to discover the causes of the abnormal market phenomena, e.g., flash crash [1, 2] and spoofing [3, 4]. Traditional machine learning methods are well-equipped to detect the anomalies but lack of the financial interpretability of the causes [5, 6]. Given observed market time series data, financial market simulation (FMS) aims to directly approximate its underlying data generating process, informed by the trading rules of the real market [7]. By simulating the underlying trading activities behind the observed market data, FMS is able to discover how the structure of the financial market changed over time and explain the changes on the micro level [8, 9].
@@ I Introduction
Advanced methods from both the fields of optimization and statistical inference have been developed to address the calibration problem [14, 15, 16]. Optimization methods primarily focus on minimizing the discrepancy through black-box optimization techniques [17], while statistical inference approaches estimate the likelihood or posterior distribution of $\hat{\mathbf{X}}_{T}$ and $\mathbf{w}$ for selected samples subject to $D(\hat{\mathbf{X}}_{T},M(\mathbf{w}))<\epsilon$ [18, 19]. In both ways, the calibration is conducted in an iterative randomized sampling manner in the parameter space and the sampling is mostly guided by the discrepancy. Intuitively, the more information observed from the market included in $\hat{\mathbf{X}}_{T}$, the closer $M(\mathbf{w})$ is expected to approximate the underlying data generative process. Most FMS works merely consider the mid-price data as $\hat{\mathbf{X}}_{T}$ [15, 20], other important information of observable data, e.g., trading volume, bid/ask directions, and inter-arrival time of orders. Consequently, the calibrated $M(\mathbf{w})$ fails to capture the underlying market structure.
Compared to mid-price data, the Limit Order Book data (LOB) provides a more detailed and informative source of market information. It offers a comprehensive view of all untraded orders, capturing critical market dynamics beyond simple price changes [21]. As the fundamental market data, LOB is widely adopted by most world-class securities exchanges [22, 23]. In LOB, all the untraded orders fall into either the ask side (for selling) or the bid side (for buying) according to their order directions. When a new order arrives, the LOB is updated immediately in one of the two ways: if the order price matches a price on the opposite side, the order is executed and the corresponding orders on the opposite side are removed; otherwise, the new order is added to its respective side. On each side, the untraded orders are organized in the ”price-first-time-second” manner [24]. Notably, the untraded orders in LOB implicitly reflect the trading intentions of the entire market. Besides, the immediate updates to the LOB provide a highly fine-grained view of market microstructure. On this basis, simulating LOB offers a conceptually close approximation to the real data generative process underlying the market scenario of interest.
@@
II Related Work
II-A Calibration of FMS
@@ II-B Deep Learning for LOB
Recent works propose to generate LOB data with deep learning, which is quite close to the goal of FMS [35]. However, they do not involve any validation or calibration steps to force the generated data resembling any given observed LOB but are only limited to following some macro properties of the market [36, 37, 38]. Besides, those deep generative models lack the interpretability of the quotes and trades in the real market. Some of the models also are not informed by the real-world matchmaking rules for updating LOB [39, 40], and all of them are not explainable in why the specific orders are submitted. These issues impose significant restrictions on the use of these models as the simulators since one can hardly intervene in the order streams to do a ”what-if” test and reveal the micro-structured causes of certain financial events [41, 42]. Furthermore, they explicitly require the ground truth order streams as input to train the networks, which is not available in the setting of FMS calibration problems.
@@
III The Proposals
III-A The Problem Definition and Challenges
@@
    Data is with different scales. Prices typically fluctuate around several thousands of cents with a tick size of 1 cent, while volumes range from 100 or 1000 shares to millions of shares, dependent on the rules of exchanges.
To summarize, learning representations of LOB requires jointly tackling the aforementioned three challenges. While the challenges of multi-scale and non-linearly auto-correlation have been extensively investigated in the literature, the precedence constraints have been rarely seen in other fields. Further more, it is uncommon to address them simultaneously with one representation model. These difficulties jointly offer LOB unique challenges and a good arena for studying representation learning of such type of data, particularly as the reconstructed multi-dimensionality data must maintain precedence constraints and avoid ”intersection” at each time step.
@@ III-B The Network Architecture
Implementation Details. The proposed autoencoder, SimLOB, processes observed LOB by spliting it into multiple segments of $\tau=100$ time steps, as suggested by [33]. Through sensitive analysis presented in Section IV-D, we set $L=2$ and $\tilde{\tau}=128$ by default. The reconstruction error is defined as:
@@
Figure 3: The workflow of the PGPS financial market simulation model.
IV Empirical Studies
To comprehensively evaluate the proposed SimLOB, several research questions (RQs) naturally arise:
RQ1: How well does SimLOB represent the latent structure of the underlying market dynamics?
RQ2: Can SimLOB generalize across different financial markets?
RQ3: Are the learned latent representation interpretable for financial applications?
RQ4: What is the recommended architecture?
To address these questions, we design four groups of experiments:
Group 1: Assessing the representation capability of SimLOB on synthetic training data, focusing on reconstruction and calibration tasks.
Group 2: Evaluating the generalization of SimLOB on real stock data and synthetic data with advanced order types and varying tick sizes, while trained on merely synthetic data.
Group 3: Investigating the interpretability of different network layers of SimLOB.
Group 4: Ablating the parameter sensitivity and structural design.
IV-A Group 1: Representation on Synthetic Data for Calibration
We conduct three main experiments in Group 1. First, the proposed SimLOB and eight state-of-the-art networks are trained on a large volume of synthetic LOB data generated by an FMS model, as introduced in Section IV-A1. The reconstruction errors on the test set are then measured to evaluate the effectiveness of the vectorized representations. Second, we show that calibrating FMSs with learned representations of LOB is more beneficial than raw LOB inputs and traditional mid-price signals. Third, latent vectors from SimLOB and 8 compared networks are applied to ten FMS calibration tasks to examine whether better LOB representations consistently lead to improved calibration performance.
TABLE I: Parameter Ranges for Generating Synthetic Data
@@ IV-A Group 1: Representation on Synthetic Data for Calibration
IV-A1 General Experimental Setup
The Adopted FMS Model. The widely studied Preis-Golke-Paul-Schneid (PGPS) model [49, 16] is adopted as the FMS model, simulating two types of agents: 125 liquidity providers submiting limit orders and 125 liquidity takers submiting market orders or cancel untraded limit orders. The simulation flowchart is shown in Fig. 3. At each time step $t$-th, a liquidity provider submits a limit order with probability $\alpha=0.5$. A liquidity taker submits a market order with probability $\mu$ and cancel an untraded limit order with probability $\delta$. The probability of the market order being either the bid side or ask side is $q_{taker}(t)$ and $1-q_{taker}(t)$, respectively, where $q_{taker}(t)$ follows a mean-reverting random walk centered at 0.5, with increments of ±$\pm\Delta s$ and reversion probability $0.5+|q_{taker}(t)-0.5|$ Each order has a fixed volume of 100 shares. The price of an ask limit order is determined by $p^{a}_{1}(t)-\lfloor-\lambda(t)\log u)\rfloor-1$ and a bid limit order is determined by $p^{b}_{1}(t)-\lfloor-\lambda(t)\log u)\rfloor+1$, where $\lambda(t)=\lambda_{0}(1+\frac{|q_{taker}(t)-0.5|}{\sqrt{\langle q_{taker}-0.5\rangle^{2}}}C_{\lambda})$. Here $\langle q_{taker}-0.5\rangle^{2}$ indicates a pre-computed value obtained by taking the average of $10^{5}$ Monte Carlo iterations of $(q_{taker}-0.5)^{2}$ before simulation. $u\sim U(0,1)$ is a uniform random number. The price of the market order is automatically set to the best price level on the opposite side. In summary, the PGPS model contains 250 agents and the key parameters to be calibrated are $\mathbf{w}=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$.
The Dataset for Representation Learning. Augmenting training data with the synthetic LOB generated by FMS is increasingly popular [35, 40]. Hence, this work trains the SimLOB with purely the synthetic LOB data generated by the PGPS model with different parameters. In general, each setting of the 6-parameter tuple of PGPS actually defines a specific market scenario. To ensure the synthetic data enjoys enough diversity, we uniformly randomly sample 2000 different settings of those 6-parameter tuples from a predefined range suggested by [29] (see Table I). Each configuration simulates over 50,000 time steps, yielding 500 sequences with 100 time steps. This results in 1 million LOB sequences ($\tau=100$), with 80% for training and 20% for testing. Additionally, real market data from the sz.000001 stock of the Chinese market (during May of 2019) is used for evaluation.
TABLE II: The 9 Compared Networks
@@ IV-A1 General Experimental Setup
The Compared Networks. This experiment compares 7 SOTA networks and one enhanced version (TransLOB-L), using their architecture (excluding the last linear layer) as representation learning layers, all configured with $\tilde{\tau}=128$. These networks have been frequently adopted in comparisons in recent deep learning based LOB analysis works [33, 50, 46], primarily focusing on price trend prediction. We follow the conventions of [33] to name the 8 networks as MLP[51], LSTM[51], CNN1[52], CNN2[53], CNN-LSTM[53], DeepLOB[46], TransLOB[44]. As their names suggest, despite the fact that MLP, LSTM, CNN1, and CNN2 adopt single-type networks, the other works mostly employ a Convolution-Recurrent architecture to intuitively first extract the prices precedence and then capture the temporal auto-correlation. TransLOB employs the Transformer architecture to replaces the recurrent network while still keeping the convolution layers. TransLOB-L is an enlarged version of TransLOB to keep the size approximately aligned with SimLOB by using 7 stacked Transformers instead of 2 as in the original TransLOB. All 8 networks are trained under the same protocol as SimLOB, with input formats and sizes follow the original papers. Table II summarizes their characteristics: the second and third columns list acceptable input formats of each network, and the last column shows the sizes of the networks. The MLP has the largest size of $2.2\times 10^{7}$ weights, which can be simply considered as the enlarged version of $\mathrm{FCN}_{2}$.
@@
TABLE III: Parameter Settings of Generating 10 Synthetic Data with the PGPS Model
@@
10 synthetic LOB data are randomly generated with 10 different settings of 6-parameter tuples, each containing $T=3600$ time steps to simulate approximately 1 hour of market activity. The parameters range for randomly sampling parameters is in the Table I, and the sampled tuples are listed in Table III. These data are utilized as the target data. Note that these data instances are challenging that they have a much higher frequency than traditional FMS works who can only calibrate to daily data [29]. And the length of the target data is also at least 10×10$\times$10 × longer than the existing calibration works [15]. The objective function for calibration employs MSE. The calibration objective for mid-prices follows traditional works do, Eq.(2) gives
@@
Figure 5: The reconstructed bid prices (colored lines represent 10 levels) on 4 synthetic data are visualized with $\mathbf{Err}_{r}$.
IV-A2 Experiment-1: Learning the Vectorized Representations of LOB
The reconstruction errors on 0.2 million synthetic testing data of all 9 networks are depicted in Fig. 4. Among them, SimLOB not only achieves the smallest averaged reconstruction error, but performs the most stably, with reconstruction errors of SimLOB following a power-law distribution and a mode of 0.003. Almost all the CNN-based networks (including CNN1, CNN2, CNN-LSTM, TransLOB, and TransLOB-L) perform the worst, suggesting that CNN is not effective as expected to deal with LOB data. This is quite contradictory to the trend of existing works on adopting deep learning for analyzing LOB. This phenomenon is later studied in-depth in Section IV-E with more empirical ablations.
The reconstructed data of seven compared networks across four synthetic data demonstrates the superiority of SimLOB. These depicted data is randomly selected from the testing instances based on the principles of covering different fluctuating amplitudes and trends.The data reconstructed by SimLOB mostly closely resembles the ground truth, while other networks have tended to smooth the target data, failing to fully preserve the key properties of LOB in their latent vectors. TransLOB and TransLOB-L are excluded from visualization as their reconstructed data appear to be nearly flat, consistent with their poor performance shown in Fig. 4. This indicates that the Transformer architecture alone is not the panacea to the representation learning of LOB. And thus it verifies that the importance of the proposed feature extraction and dimension reduction blocks in SimLOB.
Figure 6: Mid-price reconstruction on synthetic data with different levels of volatility and turnover.